OpenStudy (anonymous):

solve 4(5m + 6) - 8 ≤ 336

OpenStudy (anonymous):

Start by distributing the problem. 4(5m+6).

OpenStudy (astrophysics):

Distribute 4 over 5m+6 first\[4(5m+6)-8 \le 336\] \[20m + 24-8 \le 336\] \[20m + 16 \le 336\] now subtract 16 on both sides, then divide by 20. \[20m \le 336 - 16 \implies 20m \le 320\] \[m \le \frac{ 320 }{ 20 } \implies m \le 16\]

OpenStudy (anonymous):

thank you so much for the help

OpenStudy (astrophysics):

Np